(Assume the pully and string are massless and ther
Answer #1

Problem 7 10 kg 10g NN 5 k 5 g

Applying Newton's 2nd law,

Fnet = m a

on 5kg,

5g - T = 5a .........(i)


on 10 kg:

T = 10 a ........(ii)


(i) + (ii ) =>

5g - 10a = 5a

a = 5 x 9.81 / (5 + 10) = 3.27 m/s^2